OpenStudy (anonymous):

Find the probaility that 1 , 3 , 3 , 5 , 5 ,5 picked at random; does not end with 33

hartnn (hartnn):

a 2 digit number picked at random , right??

OpenStudy (anonymous):

a 2 digit number picked at random..? I think that you use all 6 digits

hartnn (hartnn):

ok,a 6 digit number picked at random does not end with 33....is it?

OpenStudy (anonymous):

I guess so. I tried it but i did not get the right answer.

OpenStudy (anonymous):

I got \[\frac{4!}{2!} *\frac{3*4}{3!} \]in which failed

hartnn (hartnn):

what do u get as total numbers formed? and obviously, repetition is allowed here isn't it ?

OpenStudy (anonymous):

total numbered form is obv. 60. and repetition is allowed i suppose.

hartnn (hartnn):

ok,answer my this question , how many 4 digit numbers u can form with 1,5,5,5.

OpenStudy (anonymous):

4!/3!

hartnn (hartnn):

so is the final answer 56 ?

OpenStudy (anonymous):

the answer is 56. but how

hartnn (hartnn):

i mean 56/60 is the probability. i just subtracted the total 'numbers' not valid (i.e. one with 33 in the end) from total numbers, so 60-4=56

hartnn (hartnn):

i found total numbers invalid by fixing 33 in the last 2 digits of 6 digit number. then the possible combinations are there only for 1st 4 places with 1,5,5,5. did u get it?

hartnn (hartnn):

tell me if u didn't get @Omniscience

OpenStudy (anonymous):

yes i do. \[60 - \frac{4!}{3!} *\frac{2!}{2!} = 56\] but is there by change another method; wher i dont need to minus the invalid by the total?